Crispy Chips can sell a bag of chips for $1.25 when the demand is 300 bags of chips. When the demand is 400 bags, the company can charge only $1.05 per bag. Use this information to find the demand equation assuming a linear relationship between demand (x in hundreds of bags) and price (p in dollars).
Answer #1

Let x = m*p + b

First is p = $1.25 then x = 3 (as x is in hundreds)

So, 3 = 1.25*m + b ----------(1)

Second is: p = $1.05 then x = 4

So, 4 = 1.05*m + b ----------(2)

Subtract (2) from (1): -1 = 0.2m

m = -1/0.2 = -5

Put m = -5 in equation (1):

3 = 1.25 * (-5) + b

b = 3 + 6.25 = 9.25

Hence the equation is: x = -5*p + 9.25

Please comment if any doubt. Thank you.